Two subway stops are separated by 1100 m. If a subway train accelerates at +1.2 m/s from rest through the first half of the dis tance and decelerates at -1.2 m/s through the second half, what are (a) its travel time and (b) its maximum speed? (c) Graph x, v, and a versus t for the trip.
1
Expert's answer
2020-11-16T07:48:16-0500
"s=0.5s+0.5s=0.5at^2+0.5at^2\\s=at^2=a(0.5T)^2=0.25aT^2\\1100=0.25(1.2)T^2\\T=60.6\ s"

"V^2=2a(0.5s)=as\\V^2=1.2(1100)\\V=36.3\frac{m}{s}"